OpenStudy (anonymous):

What is a quartic function with only two real zeros, x = 5 and x = 9?

OpenStudy (jdoe0001):

hmmm a quartic, meaning 4 degree so, 4 answers or solutions, or zeros so... say x = 5 x = 5 x = 9 x = 9 so there, meaning \(\begin{array}{llll} x=5\implies &x-5=0\\ x=5\implies &x-5=0\\ x=9\implies &x-9=0\\ x=9\implies &x-9=0\\ \end{array} \\ \quad \\ (x-5)(x-5)(x-9)(x-9)=\textit{original polynomial}\)

OpenStudy (phantomcrow):

A quadratic means it has a degree of 2 as in \[ax^2+bx+c\] and since x= 5 and x=9 are it's solutions, then the factored polynomial is \[(x-5)(x-9)\] and the expanded form is \[x^2-14x+45\]